Calculate the boiling point, in degrees celsius (°C), of a 4.2mole/kg aqueous solution of Al(NO3)3. Assume you are at sea level. Round your answer to three significant figures. Include the unit in your answer.
1
Expert's answer
2017-05-02T15:30:09-0400
∆t=E*m

E(H₂O)=0.52

m=4.2mol/g

∆t=0.52*4.2=2.184 C

t=100+2.184=102.184 C